OpenStudy (anonymous):

The population of a southern city follows the exponential law. If the population doubled in size over 17 months and the current population is 40,000, what will the population be 4 years from now? I've keep trying to get this right and have failed over and over. Can anyone please help me?

OpenStudy (whpalmer4):

Sure. Can you show me what you've tried?

OpenStudy (anonymous):

The problem resets each time you get it incorrect with a similar problem, do you want to see my attempts on the old problem(s) or just this one?

OpenStudy (whpalmer4):

well, you can tell me and I won't reset :-) so, with the current problem (doubling every 17 months), how would you write the general equation? Say the current population is P0...

OpenStudy (anonymous):

\[40,000e ^{K(48)} \] Where K is the rate?

OpenStudy (anonymous):

Errr A(t) = that, that is.

OpenStudy (whpalmer4):

okay, you can do it that way, but there's another way which might be easier. If we double every 17 months, then how about multiplying the population by \(2^{t/17}\) where \(t\) is in months? at t = 0 months, that will evaluate to \(2^0 = 1\). At \(t = 17\) months, that will evaluate to \(2^{17/17} = 2^1 = 2\). At \(t=34\) months, that will evaluate to \(2^{34/17} = 2^2 = 4\), etc.

OpenStudy (whpalmer4):

Saves having to compute the decay/growth coefficient (which admittedly is just saving you dividing 0.693/17)

OpenStudy (whpalmer4):

No, you don't use e at all in this form...

OpenStudy (whpalmer4):

\[P(t) = P_0 2^{t/t_{dbl}}\]where \(t_{dbl}\) is the time to double the population.

OpenStudy (whpalmer4):

or if you were doing exponential decay instead of growth, it would be the time to halve the population. the advantage of doing it this way is if you have a different sort of growth (for example, say your population triples every 17 months), you just use the different base for the exponential. tripling would be \(P_0*3^{t/17}\), etc.

OpenStudy (anonymous):

Okay, re-did it, got 283156.8

OpenStudy (anonymous):

Oh nvm that is correct, thank you. ^^

OpenStudy (whpalmer4):

Well, we can work it out both ways, and you can do whichever you like. I'll do it this way, and you do it with the decay/growth constant. \[P(t) = P_0 2^{t/17}\]\[P(t) = 40000*2^{t/17}\]but if we want to take an input value of \(t\) in years, then we need to multiply \(t*12\) to get months in our calculation, maybe this is what is tripping you up... \[P(t) = 40000*(2)^{12t/17}\]\[P(4) = 40000*(2)^{12*4/17} = 40000*2^{48/17} = 283157\]

OpenStudy (whpalmer4):

Now if we went with the e-based version, the growth factor would be \[\lambda = 12*\frac{\ln 2}{17} = 0.48928\] (including the conversion from years to months) \[P(t) = P_0 e^{\lambda t} = 40000*e^{0.48928t} \]\[P(4) = 40000*e^{0.48928*t} = 283157\]

OpenStudy (anonymous):

I understand now, thank you. :)

OpenStudy (whpalmer4):

Great! The 2^t is a bit easier for initial understanding, I think, and the e^lambda t version is easier when you start doing differential equations...